# FedGraph: A Research Library and Benchmark for Federated Graph Learning

### v3

@@ Abstract
Federated graph learning is an emerging field with significant practical challenges. While algorithms have been proposed to improve the accuracy of training graph neural networks, such as node classification on federated graphs, the system performance is often overlooked, despite it is crucial for real-world deployment. To bridge this gap, we introduce FedGraph, a research library designed for practical distributed training and comprehensive benchmarking of FGL algorithms. FedGraph supports a range of state-of-the-art graph learning methods and includes a monitoring class that evaluates system performance, with a particular focus on communication and computation costs during training. Unlike existing federated learning platforms, FedGraph natively integrates homomorphic encryption to enhance privacy preservation and supports scalable deployment across multiple physical machines with system-level performance evaluation to guide the system design of future algorithms. To enhance efficiency and privacy, we propose a low-rank communication scheme for algorithms like FedGCN that require pre-training communication, accelerating both the pre-training and training phases. Extensive experiments benchmark different FGL algorithms on three major graph learning tasks and demonstrate FedGraph as the first efficient FGL framework to support encrypted low-rank communication and scale to graphs with 100 million nodes [^1] Code available at https://github.com/FedGraph/fedgraph.
@@ 1 Introduction
Graph neural networks aim to learn representations of graph-structured data that capture features associated with graph nodes as well as edges between them (Bronstein et al., 2017). Such representations can be used to represent and classify nodes or predict the formation of links in large graphs, or can be used to represent and classify smaller graphs. Most graph applications can modeled as one of three major graph learning problems: node classification (e.g., classifying nodes representing papers in citation networks based on their research topic), link prediction (e.g., recommending the formation of links that represent friendship between users in social networks), or graph classification (e.g. classifying types of proteins in biology, where each protein is represented as a graph). Figure 1 (left) illustrates these three types of graph learning tasks (Benamira et al., 2019; Zhang et al., 2020b).
In practice, graph data is often too large to be trained on a single server or may naturally exist on multiple local clients. For example, graph learning on records of billions of users’ website visits requires significant computational resources, beyond those of a single server. Even if a single server or data center could hold such information, privacy regulations may require that it be stored where it was generated, e.g., the General Data Protection Regulation (GDPR) in Europe and Payment Aggregators and Payment Gateways (PAPG) in India prevent private user data from being shared across international borders. Users may also not want to share their personal data with an external server. In response to these concerns, federated learning has been widely studied as a way to preserve user privacy while training accurate models on data stored at multiple clients (Zhao et al., 2018).
@@
These challenges to training on real-world graphs motivate Federated Graph Learning (FGL) as an important research topic (Liu et al., 2024). The training schematic of such a federated architecture is shown in Figure 1 (right): each client maintains a local graph and a local Graph Neural Network (GNN) model, which can be aggregated with other clients’ local models at a coordinator server. Within this architecture, there are many different FGL algorithms that handle local model updates and aggregation differently, resulting in varying model accuracies and system performance (e.g., runtime, communication costs) (Xie et al., 2021; Zhang et al., 2021; Yao et al., 2024). While evaluating the model accuracy of such algorithms can be easily done with simulations on open graph datasets (Hu et al., 2020), evaluating their real-world system performance requires sophisticated benchmarking platforms that can fairly compare multiple FGL algorithms.
Vanilla FGL benchmark platforms are generally designed for research purposes only and simulate the actions of multiple clients on a single machine (Xie et al., 2021; Li et al., 2024). Thus, they do not capture the real communication and computation costs of training graph models in a federated manner across distinct physical clients. Meanwhile, existing federated learning libraries mainly focus on overcoming client data heterogeneity, with limited support for GNN training. For example, FedScale (Lai et al., 2022) does not support graph models, while FedGraphNN (He et al., 2021) and FederatedScope-GNN (Wang et al., 2022) are not well-maintained and only support basic federated learning methods like FedAvg. Users need to implement the FGL algorithms themselves, including steps unique to FGL tasks such as handling cross-client graph edges that begin and end at different clients. Moreover, none of these platforms natively support FGL enhancements that may be needed in practice, e.g., homomorphic encryption for enhanced privacy (Jin et al., 2023; Yao et al., 2024).
@@
  • Extensive experiments benchmark the system performance of various algorithms on three major FGL tasks and showcase the ability to support a privacy-preserving federated system on graphs with 100 million nodes.
@@
Table 1: FedGraph Comparison with Existing Frameworks. FedGraph supports distributed training with various FGL algorithms, including algorithms that specifically support cross-client edges and encrypted aggregation. With a system-level profiler, FedGraph can then benchmark existing algorithms and provide optimizations to support training on large (tens of millions of nodes) graphs.
In this paper, we begin with an overview of the system design in Section 2, followed by highlighting the key system components in Section 3. In Section 4, we present a case study demonstrating how FedGraph facilitates the easy design and test of low-rank pre-training communication in FGL. We then benchmark the performance on three major tasks and evaluate its scalability in Sections 5. We finally conclude the paper in Section 6.
@@ 2.1 Design Principles
Optimized usability with simple configurations for simulation and federated training: At the access layer, users configure training settings and can seamlessly switch between local simulation and federated training using the same codebase. With just 10–20 lines of code, users can train a federated GNN model, as we discuss further in Section 2.2. FedGraph abstracts away the complexity of federated graph training, offering a unified platform for training and evaluation.
Benchmark existing federated graph learning methods: In the application layer, FedGraph supports three FGL tasks (node classification, link prediction, and graph classification) and a wide range of state-of-the-art model training algorithms for each task, implemented either by the original authors or library developers. Appendix C lists the currently supported algorithms and datasets.
@@ 2.2 FedGraph Use Example
Researchers can easily install the FedGraph library by running pip install fedgraph in Python, configure experiments, and start training federated GNN models. Appendices AB, and D provide more details about the code structure and user API. As illustrated in Figure 2 (right), users can initiate a federated graph learning experiment through a straightforward configuration process. Task, method, and dataset specification: The user specifies the learning task, algorithm, dataset, and client data distribution. FedGraph enforces explicit task-method combinations to ensure compatibility and reproducibility. Training configuration: Key hyperparameters, such as the number of global rounds, local training steps, learning rate, and number of participating trainers, are defined in the configuration. Security and privacy settings: Users can enable homomorphic encryption to ensure privacy-preserving training. Execution: Once settled configuration, calling the API function run_fedgraph(config) launches the experiment. FedGraph then automatically handles data loading, client initialization, and distributed training across clients.
@@ 3.1 FedGraph Monitoring System
The FedGraph Monitoring System, illustrated in Figure 3 (right), uses a Monitor Class to track key system metrics, including running time, CPU/GPU utilization, memory consumption, and communication costs between the server and clients. Examples of the monitoring dashboard during training are provided in Appendix E.
@@ 3.2 FedGraph Homomorphic Encryption
FedGraph supports homomorphic encryption (HE) to enable end-to-end secure computation, ensuring user privacy during training (Figure 4). To our knowledge, it is the first FGL library to incorporate end-to-end HE as a core capability. While prior work has applied HE to specific FGL scenarios (Yao et al., 2024), FedGraph generalizes this support by addressing two key requirements of realistic FGL settings: (i) secure feature aggregation prior to training (Yao et al., 2024), and (ii) secure model aggregation during training (Zhang et al., 2024; Kim et al., 2025).
@@
Federated Aggregation During Training: During model aggregation, FedGraph encrypts each client’s model updates before transmission. The central server then performs homomorphic aggregation on the encrypted updates to compute the global model without accessing any plaintext data (Figure 4(Zhang et al., 2020a; Jin et al., 2023).
@@ 3.3 FedGraph Kubernetes
To support diverse training workloads and scale to a large number of clients, FedGraph employs Kubernetes as the foundation of its infrastructure layer. Unlike prior platforms (Table 1) that lack native Kubernetes integration, FedGraph is designed to scale efficiently to both large client populations and massive graph datasets. As shown in Figure 6, FedGraph leverages Amazon Web Services Elastic Kubernetes Service to enable flexible and dynamic resource management. A self-managed Kubernetes cluster equipped with GPU instances is deployed to handle computationally intensive tasks. The architecture consists of a master node responsible for orchestration, task scheduling, resource allocation, and cluster health monitoring, while worker nodes execute distributed training processes in parallel. To enhance scalability and cost-efficiency, we integrate the Kubernetes Cluster Autoscaler, which dynamically adjusts worker nodes based on workload demands. This ensures resources are used only when needed, maintaining high efficiency under varying system loads for continuous large-scale execution.
@@ 3.4 Optimizations for Scalability
Client Selection for Large-Scale Federated Learning: FedGraph uses a selective client engagement mechanism, allowing the user to specify a fraction of clients that send the model update at each training round (Li et al., 2019). This approach reduces communication costs at the server and resource usage at the clients, enabling scalability to thousands of clients.
@@ 4.2 Low Rank Method for Pre-Train Feature Aggregation
In FedGCN, pre-training communication involves aggregating feature information across clients for nodes with cross-client edges, which may be encrypted and securely aggregated in order to preserve privacy (Yao et al., 2024). We implement a low-rank method for pre-train feature aggregation using client-side projection of feature information to reduce communication cost. The server first generates a random projection matrix $\mathbf{P}\in\mathbb{R}^{d\times k}$ where $k\ll d$ is the specified rank (e.g., $k=100$). The server then distributes $\mathbf{P}$ to all clients. Each client $i$ computes the projected feature matrix $\mathbf{\hat{X}}_{i}=\mathbf{X}_{i}\mathbf{P}$ where $\mathbf{\hat{X}}_{i}\in\mathbb{R}^{n_{i}\times k}$ with $n_{i}$ nodes and then sends $\mathbf{\hat{X}}_{i}$ to the server. The server aggregates these low-rank projections $\mathbf{\hat{X}}_{agg}=\sum_{i=1}^{m}\mathbf{\hat{X}}_{i}$, where $m$ is the number of clients, and distributes the result back to clients. To prevent potential reconstruction of another client’s original features, the projection matrix $\mathbf{P}$ can also be encrypted before being distributed. This adds a layer of protection against inversion attacks on the shared aggregated features. This approach significantly reduces communication overhead in both directions due to only communicating low-rank information. Since our HE interface supports addition on encrypted data, it can safely perform the aggregation procedure on encrypted projected features, preserving client privacy.
@@ 4.3 Performance Evaluation
We evaluate the impact of low-rank compression on communication cost, training time, and accuracy in FedGCN using the Cora dataset. Compression ranks range from the full 1433-dimensional features to rank 100, yielding up to 93% reduction. As shown in Figure 7, HE alone introduces significant communication overhead, especially in pre-training, but this is substantially mitigated by applying low-rank projection. Across both plaintext and HE settings, accuracy remains stable even as rank decreases, demonstrating the effectiveness of low-rank compression for communication-efficient and privacy-preserving FGL.
@@ 5 Benchmarking FedGraph on Graph Learning Tasks and Scalability
5.1 Federated Graph Classification
@@
The graph classification experiment evaluates the performance of various federated learning algorithms on five benchmark datasets: IMDB-BINARY, IMDB-MULTI, MUTAG, BZR, and COX2. These datasets span diverse graph structures, including social and molecular networks, providing a comprehensive testbed for assessing algorithm effectiveness.
As shown in Figure 8, the evaluation spans 200 training rounds across five datasets. GCFL+ and GCFL+dWs consistently achieve the highest accuracy, particularly on BZR and COX2. However, as illustrated in the middle and right plots, these gains come at the cost of significantly higher training time and communication overhead, especially on complex datasets like IMDB, reflecting their greater computational and communication demands. In contrast, FedAvg offers the lowest communication cost and the shortest, most consistent training time across all datasets, making it a practical choice for latency- and bandwidth-constrained environments.
5.2 Federated Node Classification
@@
We evaluate FedAvg and FedGCN on four node classification benchmarks: Cora, Citeseer, PubMed, and Ogbn-Arxiv. As shown in Figure 9 (left), FedGCN consistently achieves higher accuracy than FedAvg across all datasets. However, this improvement comes at the cost of significantly longer training times, particularly on larger datasets such as Ogbn-Arxiv in Figure 9 (middle).
The communication cost breakdown (Figure 9, right) reveals the source of this overhead: FedGCN incurs substantially higher pre-training communication costs compared to FedAvg. Notably, the observed communication costs closely match the theoretical values, validating the efficiency modeling of these algorithms. This is especially pronounced for Ogbn-Arxiv, where the pre-training phase dominates the total communication volume. While FedAvg has relatively balanced communication between pre-training and training phases, FedGCN’s architecture leads to a disproportionate communication burden during the pre-training aggregation of features across clients. This imbalance helps explain the increased total training time observed for FedGCN on large-scale datasets.
5.3 Federated Link Prediction
@@
In the federated link prediction setting, we simulate a scenario where each client holds region-specific data from the Foursquare Global-scale Check-in Dataset (Yang et al., 2016), covering three geographic configurations: US; US and BR; and US, BR, ID, TR, and JP. This setup respects data privacy by avoiding raw data sharing across regions. We evaluate four algorithms, 4D-FED-GNN+, FedLink, STFL, and StaticGNN, across three key metrics: AUC for predictive accuracy, training time for computational efficiency, and communication cost for network efficiency.
@@
Figure 11: Accuracy Across Global Training Rounds and CPU Usage for Node Classification Tasks.

### v4

@@ Abstract
Federated graph learning is an emerging field with significant practical challenges. While algorithms have been proposed to improve the accuracy of training graph neural networks, such as node classification on federated graphs, the system performance is often overlooked, despite it is crucial for real-world deployment. To bridge this gap, we introduce FedGraph, a research library designed for practical distributed training and comprehensive benchmarking of FGL algorithms. FedGraph supports a range of state-of-the-art graph learning methods and includes a monitoring class that evaluates system performance, with a particular focus on communication and computation costs during training. Unlike existing federated learning platforms, FedGraph natively integrates homomorphic encryption to enhance privacy preservation and supports scalable deployment across multiple physical machines with system-level performance evaluation to guide the system design of future algorithms. To enhance efficiency and privacy, we propose a low-rank communication scheme for algorithms like FedGCN that require pre-training communication, accelerating both the pre-training and training phases. Extensive experiments benchmark FGL algorithms on three major graph learning tasks and demonstrate FedGraph as the first efficient FGL framework to support encrypted low-rank communication and scale to graphs with 100 million nodes.
@@ 1 Introduction
Graph neural networks aim to learn representations of graph-structured data that capture features associated with graph nodes as well as edges between them [1]. Most graph applications can modeled as one of three major graph learning problems: node classification (e.g., classifying nodes representing papers in citation networks based on their research topic), link prediction (e.g., recommending the formation of links that represent friendship between users in social networks), or graph classification (e.g. classifying types of proteins in biology, where each protein is represented as a graph). Figure 1 (left) illustrates these three types of graph learning tasks [2, 3].
In practice, graph data is often too large to be trained on a single server or may naturally exist on multiple local clients. For example, graph learning on records of billions of users’ website visits requires significant computational resources, beyond those of a single server. Even if a single server or data center could hold such information, privacy regulations may require that it be stored where it was generated, e.g., the General Data Protection Regulation (GDPR) in Europe and Payment Aggregators and Payment Gateways (PAPG) in India prevent private user data from being shared across international borders. Users may also not want to share their personal data with an external server. In response to these concerns, federated learning has been widely studied as a way to preserve user privacy while training accurate models on data stored at multiple clients [4].
@@
These challenges to training on real-world graphs motivate Federated Graph Learning (FGL) as an important research topic [5]. The training schematic of such a federated architecture is shown in Figure 1 (right): each client maintains a local graph and a local Graph Neural Network (GNN) model, which can be aggregated with other clients’ local models at a coordinator server. Within this architecture, there are many different FGL algorithms that handle local model updates and aggregation differently, resulting in varying model accuracies and system performance (e.g., runtime, communication costs) [6, 7, 8]. While evaluating the model accuracy of such algorithms can be easily done with simulations on open graph datasets [9], evaluating their real-world system performance requires sophisticated benchmarking platforms that can fairly compare multiple FGL algorithms.
Vanilla FGL benchmark platforms are designed for research purposes only and simulate the actions of multiple clients on a single machine [6, 10]. Thus, they do not capture the real communication and computation costs of training graph models in a federated manner across distinct physical clients. Meanwhile, existing libraries mainly focus on overcoming client data heterogeneity, with limited support for GNN training. For example, FedScale [11] does not support graph models, while FedGraphNN [12] and FederatedScope-GNN [13] are not well-maintained and only support basic federated learning methods like FedAvg. Users need to implement the FGL algorithms themselves, including steps unique to FGL tasks such as handling cross-client graph edges that begin and end at different clients. Moreover, none of these platforms natively support FGL enhancements that may be needed in practice, e.g., homomorphic encryption for enhanced privacy [14, 8].
@@
  • Extensive experiments benchmark system performance on three major FGL tasks and showcase support for a privacy-preserving federated system on graphs with 100 million nodes.
Table 1: Comparison with Existing Frameworks. FedGraph supports distributed FGL, cross-client edges, encrypted aggregation, and system-level profiling for large-scale optimization.
@@
In this paper, we first overview the system design in Section 2, followed by highlighting the key system components in Section 3. In Section 4, we present a case study demonstrating how FedGraph facilitates the design and test of low-rank pre-training communication in FGL. We then benchmark the performance on three tasks and evaluate its scalability in Sections 5, and conclude in Section 6.
@@ 2.1 Design Principles
Optimized usability with simple configurations for simulation and federated training: At the access layer, users configure training settings and can seamlessly switch between local simulation and federated training using the same codebase. With just 10-20 lines of code, users can train a federated GNN model, as we discuss further in Section 2.2. FedGraph abstracts away the complexity of federated graph training, offering a unified platform for training and evaluation.
Benchmark existing federated graph learning methods: In the application layer, FedGraph supports three FGL tasks (node classification, link prediction, and graph classification) and a wide range of state-of-the-art model training algorithms for each task, implemented either by the original authors or library developers. Appendix D lists the currently supported algorithms and datasets.
@@ 2.2 FedGraph Use Example
Researchers can easily install the FedGraph library by running pip install fedgraph in Python, configure experiments, and start training federated GNN models. Appendices BC, and E provide more details about the code structure and user API. As illustrated in Figure 2 (right), users can initiate a federated graph learning experiment through a straightforward configuration process. Task, method, and dataset specification: The user specifies the learning task, algorithm, dataset, and client data distribution. FedGraph enforces explicit task-method combinations to ensure compatibility and reproducibility. Training configuration: Key hyperparameters, such as the number of global rounds, local training steps, learning rate, and number of participating trainers, are defined in the configuration. Security and privacy settings: Users can enable homomorphic encryption to ensure privacy-preserving training. Execution: Once settled configuration, calling the API function run_fedgraph(config) launches the experiment. FedGraph then automatically handles data loading, client initialization, and distributed training across clients.
@@ 3.1 FedGraph Monitoring System
The FedGraph Monitoring System, illustrated in Figure 3 (right), uses a Monitor Class to track key system metrics, including running time, CPU/GPU utilization, memory consumption, and communication costs between the server and clients. Examples of the monitoring dashboard are provided in Figure 11.
@@ 3.2 FedGraph Homomorphic Encryption
FedGraph supports homomorphic encryption (HE) to enable end-to-end secure computation, ensuring user privacy during training (Figure 4). To our knowledge, it is the first FGL library to incorporate HE as a core capability. While prior work has applied HE to specific FGL scenarios [8], FedGraph generalizes this support by addressing two key requirements of realistic FGL settings: (i) secure feature aggregation prior to training [8], and (ii) secure model aggregation during training [15, 16].
@@
Federated Aggregation During Training: Each client first encrypts its model update before transmission. The central server then performs homomorphic aggregation on the encrypted updates to compute the global model without accessing any plaintext data (Figure 4[17, 14].
@@ 3.3 FedGraph Kubernetes
To support diverse training workloads and scale to a large number of clients, FedGraph employs Kubernetes as the foundation of its infrastructure layer. Unlike prior platforms (Table 1) that lack native Kubernetes integration, FedGraph is designed to scale efficiently to both large client populations and massive graph datasets. As shown in Figure 6, FedGraph leverages Amazon Web Services Elastic Kubernetes Service to enable flexible and dynamic resource management. A self-managed Kubernetes cluster equipped with GPU instances is deployed to handle computationally intensive tasks. The architecture consists of a master node responsible for orchestration, task scheduling, resource allocation, and cluster health monitoring, while worker nodes execute distributed training processes in parallel. To enhance scalability and cost-efficiency, we integrate the Kubernetes Cluster Autoscaler, which dynamically adjusts worker nodes based on workload demands. This ensures resources are used only when needed, maintaining high efficiency under varying system loads.
@@ 3.4 Optimizations for Scalability
Client Selection for Large-Scale Federated Learning: FedGraph uses a selective client engagement mechanism in Appendix A.1, allowing the user to specify a fraction of clients that send the model update at each training round [18]. This approach reduces communication costs at the server and resource usage at the clients, enabling scalability to thousands of clients.
@@ 4.2 Low Rank Method for Pre-Train Feature Aggregation
In FedGCN, pre-training communication involves aggregating feature information across clients for nodes with cross-client edges, which may be encrypted and securely aggregated in order to preserve privacy [8]. We implement a low-rank method for pre-train feature aggregation using client-side projection of feature information to reduce communication cost. The server first generates a random projection matrix $\mathbf{P}\in\mathbb{R}^{d\times k}$ where $k\ll d$ is the specified rank (e.g., $k=100$). The server then distributes $\mathbf{P}$ to all clients. Each client $i$ computes the projected feature matrix $\mathbf{\hat{X}}_{i}=\mathbf{X}_{i}\mathbf{P}$ where $\mathbf{\hat{X}}_{i}\in\mathbb{R}^{n_{i}\times k}$ with $n_{i}$ nodes and then sends $\mathbf{\hat{X}}_{i}$ to the server. The server aggregates these low-rank projections $\mathbf{\hat{X}}_{agg}=\sum_{i=1}^{m}\mathbf{\hat{X}}_{i}$, where $m$ is the number of clients, and distributes the result back to clients. To prevent potential reconstruction of another client’s original features, the projection matrix $\mathbf{P}$ can also be encrypted before being distributed. This adds a layer of protection against inversion attacks on the shared aggregated features. This approach significantly reduces communication overhead in both directions due to only communicating low-rank information. Since our HE interface supports addition on encrypted data, it can safely perform the aggregation procedure on encrypted projected features, preserving client privacy.
@@ 4.3 Performance Evaluation
We evaluate the impact of low-rank compression on communication cost, training time, and accuracy in FedGCN using the Cora dataset. Compression ranks range from the full 1433-dimensional features to rank 100, yielding up to 93% reduction. As shown in Figure 7, HE alone introduces high communication overhead, especially in pre-training, but this is substantially mitigated by applying low-rank projection. Accuracy remains stable even as rank decreases, demonstrating the effectiveness of low-rank compression for communication-efficient and privacy-preserving FGL.
@@ 5 Benchmarking FedGraph on Graph Learning Tasks and Scalability
5.1 Benchmarking Federated Graph Learning Tasks
We first benchmark FedGraph across three representative graph learning tasks, evaluating accuracy, training time, and communication cost to provide a view of the system and model performance.
5.1.1 Federated Graph Classification
@@
The graph classification experiment evaluates the performance of various federated learning algorithms on five benchmark datasets: IMDB-BINARY, IMDB-MULTI, MUTAG, BZR, and COX2. These datasets span diverse graph structures, including social and molecular networks, providing a comprehensive testbed for assessing algorithm effectiveness. As shown in Figure 8, the evaluation spans 200 training rounds across five datasets. GCFL+ and GCFL+dWs consistently achieve the highest accuracy, particularly on BZR and COX2. However, as illustrated in the middle and right plots, these gains come at the cost of significantly higher training time and communication overhead, especially on complex datasets like IMDB, reflecting their greater computational and communication demands. In contrast, FedAvg offers the lowest communication cost and the shortest, most consistent training time across all datasets, making it a practical choice for latency- and bandwidth-constrained environments.
5.1.2 Federated Node Classification
@@
We evaluate FedAvg and FedGCN on four node classification benchmarks: Cora, Citeseer, and PubMed. As shown in Figure 9 (left), FedGCN consistently achieves higher accuracy than FedAvg across all datasets. The communication cost breakdown (Figure 9, right) reveals the source of this overhead: Compared to FedAvg, which does not have a pre-training round, FedGCN still has high pre-training communication costs due to aggregation of feature sums across clients. Notably, the observed communication costs closely match the theoretical values, validating the efficiency modeling of these algorithms. Such observation inspires the design of the new low rank algorithm in Section 4.
5.1.3 Federated Link Prediction
@@
In the federated link prediction setting, we simulate a scenario where each client holds region-specific data from the Foursquare Global-scale Check-in Dataset [19], covering three geographic configurations: US; US and BR; and US, BR, ID, TR, and JP. This setup respects data privacy by avoiding raw data sharing across regions. We evaluate four algorithms, 4D-FED-GNN+, FedLink, STFL, and StaticGNN, across three key metrics: AUC for predictive accuracy, training time for computational efficiency, and communication cost for network efficiency.